### v1

# Synthetic Data Generation for Culturally Nuanced Commonsense Reasoning in Low-Resource Languages

Abstract
Quantifying reasoning capability in low-resource languages remains a challenge in NLP due to data scarcity and limited access to annotators. While LLM-assisted dataset construction has proven useful for medium- and high-resource languages, its effectiveness in low-resource languages, particularly for commonsense reasoning, is still unclear. In this paper, we compare three dataset creation strategies: (1) LLM-assisted dataset generation, (2) machine translation, and (3) human-written data by native speakers, to build a culturally nuanced story comprehension dataset. We focus on Javanese and Sundanese, two major local languages in Indonesia, and evaluate the effectiveness of open-weight and closed-weight LLMs in assisting dataset creation through extensive manual validation. To assess the utility of synthetic data, we fine-tune language models on classification and generation tasks using this data and evaluate performance on a human-written test set. Our findings indicate that LLM-assisted data creation outperforms machine translation.
Synthetic Data Generation for Culturally Nuanced Commonsense Reasoning in Low-Resource Languages
Salsabila Zahirah Pranidathanks: Equal contribution  Rifo Ahmad Genadi11footnotemark: 1  Fajri Koto Department of Natural Language Processing, MBZUAI {salsabila.pranida,rifo.genadi,fajri.koto}@mbzuai.ac.ae
Figure 1: Proposed pipeline, including the train set overgeneration process
1 Introduction
Reasoning—the ability to draw conclusions, make inferences, and understand relationships between concepts based on available information—is a key capability assessed in recent developments of large language models (Dubey et al., 2024; OpenAI et al., 2024; Hurst et al., 2024; Almazrouei et al., 2023). It has been widely studied in English through various benchmarks, including StoryCloze Mostafazadeh et al. (2016, 2017), WinoGrande Sakaguchi et al. (2021), and HellaSwag Zellers et al. (2019), which evaluate different aspects of reasoning such as story comprehension, pronoun resolution, and natural language inference. Beyond English, research on reasoning has largely relied on machine translation Ponti et al. (2020); Lin et al. (2022), often overlooking the cultural context of the target languages. While some datasets have been developed from scratch for medium-resource languages like Indonesian Koto et al. (2024), reasoning in low-resource languages remains significantly underexplored. The primary challenges include the high cost of annotation, difficulties in recruiting native speakers, and the limited availability of culturally relevant text.
In this paper, we ask: to what extent can large language models (LLMs) assist in creating reasoning datasets for low-resource languages? How do they compare to machine translation and building datasets from scratch? While LLM-assisted dataset construction has been widely explored, most efforts have focused on English Liu et al. (2022, 2024a) and traditional NLP tasks such as named entity recognition (Mohit, 2014) and question answering Ghazaryan et al. (2025); Putri et al. (2024). However, their potential for generating reasoning datasets in low-resource languages remains under-studied, particularly for culturally nuanced story comprehension tasks.
Unlike sentence-pair classification tasks for assessing reasoning (e.g., the COPA dataset), we take a story comprehension approach to evaluate LLMs’ ability to generate coherent narratives in Javanese and Sundanese over longer texts (five sentences). Specifically, we follow the StoryCloze framework Mostafazadeh et al. (2016, 2017) to assess commonsense reasoning in daily life scenarios, where a four-sentence premise is provided along with two possible options for the concluding sentence (see Figure 1).
We focus on Javanese and Sundanese because, despite having over 80 million and 32 million speakers, respectively Aji et al. (2022), these languages remain underrepresented in language technology development. Beyond their linguistic diversity, both languages are deeply intertwined with distinct cultural traditions that influence their usage. Sundanese, for instance, conveys social nuanced through varying degrees of politeness and phonetic distinctions, while Javanese employs a structuted system of speech levels to reflect hierarchical relationships in communication (Wolff and Poedjosoedarmo, 1982).
Our contribution can be summarized as follows:
  • We benchmark three open-weight and three closed-weight LLMs for StoryCloze-style synthetic data generation in Javanese and Sundanese through extensive human evaluation.
  • We introduce the first commonsense reasoning datasets for Javanese and Sundanese. For training data, we use two strategies: (1) LLM-generated data with few-shot prompting and (2) machine translation. For the test set, we employ two approaches: (1) machine translation with post-editing and (2) manually written samples by native speakers based on predefined topics, with extensive validation to ensure quality. In total, we release 1.22K high-quality LLM-generated samples, 1K human-reviewed machine-translated texts, and 1.12K human-written samples [^1] Our data are available at https://huggingface.co/datasets/rifoag/javanese_sundanese_story_cloze.
  • We conducted extensive experiments in both classification and generation settings to assess the utility of LLM-generated data and machine translation.
Javanese Sundanese
Original
Premise: Ing Yogyakarta, gudeg dadi panganan khas. Gudeg dimasak suwene jam-jaman nganggo nangka enom lan santan. Biasane disajikake karo krecek, endog, lan ayam suwir. Wong Yogyakarta biasane mangan gudeg ing esuk utawa nalika ana acara khusus. Correct: Gudeg diwarisake turun-temurun lan dadi bagian saka budaya kuliner Yogyakarta. Incorrect: Gudeg biasane digoreng nganggo lenga panas lan disajikake pedhes. Premise: Mang Tarya indit ka sawah keur ngarambét. Geus loba jukut nu mucunghul ngaganggu paré. ngarambét ngabaladah biasana dilaksanakeun tilu puluh poé saprak tandur. Mun ngarambét mindo dilaksanakeun lima puluh poe saprak tandur. Correct: Ngarambét dilaksanakeun pikeun ngaberesihan jukut di petakan sawah. Incorrect: Ngarambét dilaksanakeun saméméh panén.
English translation
Premise: In Yogyakarta, gudeg is a traditional dish. It is cooked for hours using young jackfruit and coconut milk. It is usually served with krecek, eggs, and shredded chicken. People in Yogyakarta typically eat gudeg in the morning or during special occasions. Correct: Gudeg is passed down through generations and has become a part of Yogyakarta’s culinary heritage. Incorrect: Gudeg is usually deep-fried in hot oil and served spicy. Premise: Mang Tarya went to the rice field to do ngarambét. There are already lots of wild grass that may harm the rice crops. Ngarambét ngabaladah is done thirty days after tandur. Meanwhile, Ngarambét mindo is done fifty days after tandur. Correct: Ngarambét is done to clean up wild grass in the rice fields. Incorrect: Ngarambét is done before harvest.
Table 1: Examples of story cloze in Javanese and Sundanese, with English translations for illustrative purposes.
2 Related Works
LLM-Generated Data Creation
One possible solution to tackle data scarcity in NLP is applying data augmentation (Feng et al., 2021; Ding et al., 2020; Ahmed and Buys, 2024; Liu et al., 2024b; Yong et al., 2024; Guo and Chen, 2024; Liu et al., 2022). These methods leverage the generative capabilities of LLMs to produce high-quality synthetic data that can sometimes complement or even substitute manual dataset creation.
Prior studies have explored techniques for enhancing synthetic data generation, primarily for classification tasks. WANLI (Liu et al., 2022) used GPT-3 (Brown et al., 2020) to generate synthetic natural language inference data in English Bowman et al. (2015), followed by human refinement. Yong et al. (2024) leveraged LLMs to generate synthetic English data for sentiment analysis and topic classification, then applied a bilingual lexicon for word-to-word translation into low-resource languages. In contemporaneous work, Putri et al. (2024) utilized GPT-4 OpenAI et al. (2024) for question-answering tasks in one low-resource language. Unlike these works, we focus specifically on reasoning tasks in two low-resource languages and compare multiple strategies, including LLM-assisted generation using both open-weight and closed-weight models, machine translation, and human-written data by native speakers. To ensure topic coverage, we provide predefined topics and example seeds for both LLM-assisted generation and human writing.
English Commonsense Reasoning
Story comprehension has long been a significant challenge in NLP due to the complex interplay of causal and temporal relationships within narratives. A landmark contribution to the evaluation of these aspects is the Story Cloze test, introduced by Mostafazadeh et al. (2016, 2017). Another set of benchmarks for assessing commonsense reasoning includes WinoGrande Sakaguchi et al. (2021), COPA (Gordon et al., 2012), and HellaSwag (Zellers et al., 2019).
WinoGrande (Sakaguchi et al., 2021) is designed to measure the resolution of pronouns in ambiguous contexts. The Choice of Plausible Alternatives (COPA) (Gordon et al., 2012) evaluates causal reasoning by presenting two possible continuations of a given premise, requiring models to infer the most plausible outcome. HellaSwag (Zellers et al., 2019) builds on the SWAG Zellers et al. (2018) dataset by increasing difficulty through adversarial filtering, which ensures that the correct continuation of a given context is not easily identified by simple statistical cues, but rather requires deep reasoning. These benchmarks collectively contribute to a more comprehensive evaluation of commonsense understanding in NLP models.
Commonsense Reasoning in Languages Beyond English
Evaluating commonsense reasoning across multiple languages has gained attention through multilingual benchmarks such as XCOPA (Ponti et al., 2020) and XStoryCloze (Lin et al., 2022). These benchmarks extend existing commonsense reasoning datasets to a wider set of languages by translating English corpora into various target languages. However, such translation-based approaches often fail to capture cultural nuances and may introduce artifacts that affect naturalness and contextual relevance. Additionally, neither XCOPA nor XStoryCloze include benchmarks for Sundanese or Javanese, leaving a gap in evaluating commonsense reasoning in these widely spoken Indonesian local languages.
Several high-quality Indonesian commonsense reasoning datasets have been developed, primarily through human annotation, including IndoCloze Koto et al. (2022a) for narrative understanding, COPAL-ID Wibowo et al. (2024) integrating Indonesian cultural knowledge (i.e., Jakarta region), and IndoCulture Koto et al. (2024) for cultural reasoning. While these resources support commonsense reasoning in Indonesian, a medium-resource language, extensive datasets for local languages remain scarce. In contrast, our work focuses on low-resource languages, specifically Javanese and Sundanese, two of Indonesia’s most widely spoken local languages. We build on IndoCloze as a reference for synthetic data generation and complement it with manually written stories. Unlike prior datasets that rely on MCQs or sentence-pair tasks, we emphasize full story comprehension. To ensure diversity, we adopt multiple data creation approaches, including LLM-generated, machine-translated, and human-written narratives, enabling a broader and more culturally rich evaluation setting.
3 Dataset Construction
As shown in Figure 1, our dataset construction is describued in two streams: training and testing—using IndoCloze Koto et al. (2022a) as an initial source. IndoCloze is the only story comprehension dataset for commonsense reasoning evaluation in Indonesian and incorporates Indonesian cultural elements. Each sample consists of a four-sentence premise followed by two candidate endings: one correct and one incorrect, representing the fifth sentence.
We use Indonesian dataset in our data construction as it is the national language and shares cultural elements with Javanese and Sundanese, making it a relevant source of seed data. In addition to leveraging IndoCloze, we also create a dataset from scratch by asking native speakers to write and validate samples to ensure cultural and linguistic authenticity. Similar to Mostafazadeh et al. (2016) and Koto et al. (2022a), our dataset focuses on everyday events within Javanese and Sundanese cultural contexts, incorporating relevant locations, names, foods, and cultural practices.
3.1 Training Data
Our training set is constructed using two strategies: (1) LLM-assisted data generation and (2) machine translation.
LLM-assisted Data Generation
To assess the capability of different LLMs in generating meaningful training data, we create synthetic datasets using three open-weight models: Mixtral-8x7B-Instruct (Jiang et al., 2024), Gemma2-27B-it (Rivière et al., 2024), and Llama3.1-70B (Dubey et al., 2024)—and three closed-weight models: GPT-4o (Hurst et al., 2024; OpenAI et al., 2024), Cohere Command-R-Plus (Cohere, 2024), and Claude-3-Opus (Anthropic, 2024).
For each language and LLM, we provide seed examples and predefined topics. To create seed samples, we manually translate 50 samples from IndoCloze’s training set into Javanese and Sundanese, ensuring cultural localization. For example, we adapt references to food, such as Gudeg (a traditional breakfast from Yogyakarta), and traditions like Sawéran (a ritual in Sundanese wedding ceremonies).
To ensure that the generated examples align with the intended cultural context, we use cultural topics derived from IndoCulture (Koto et al., 2024), covering areas such as food, weddings, family relationships, pregnancy and children, death, religious holidays, agriculture, fisheries and trade, art, traditional games, daily activities, and socio-religious aspects of life. The full prompt used in this study is provided in Figure 10.
Each LLM independently generate 2,000 samples (1,000 per language), resulting in six distinct datasets, each treated as a separate training set. For each LLM call, we randomly select five seed samples and prompted the model to generate new examples. The temperature is set to 0.7 to introduce variability in the generated outputs. The generation process is repeated until each LLM produced at least 1,000 valid examples. On average, each model generated approximately 166 samples per topic (see Appendix D for topic distribution).
Since not all LLM-generated samples are high quality, we train an XLM-R binary classifier to distinguish between good and bad samples. To create training data, we ask workers to rate the quality of 600 LLM-generated outputs (details provided in Section 4.2). These rated samples are then used to train the classifier. [^2] The classifier achieves 68.33% accuracy on the dev set (20% of the train set), highlighting the complexity of the task. The experimental setup for the data filtering classifier is described in Appendix A. Once trained, we apply the classifier to all LLM-generated data. After filtering, the dataset is reduced from 12,000 generated samples to 592 Sundanese and 628 Javanese high-quality samples. About half of the data generated by Claude and GPT-4o are retained, while a large portion of samples from other LLMs, especially Mixtral, are filtered out. More details on the filtered data proportions can be found in Appendix C.
Machine Translation
To compare LLM-assisted dataset construction with machine translation, we translate 1,000 instances from the original IndoCulture training set into 1,000 Javanese and 1,000 Sundanese examples. Both translations are generated using Google Translate [^3] https://translate.google.com/, accessed in September 2024., chosen for its strong translation performance in both languages. Since this dataset is intended for training, no human validation was conducted. From this point forward, we refer to it as MTtrain.
3.2 Test Set
We carefully construct the test set using two strategies: (1) machine translation with human verification, resulting in 500 Javanese and 500 Sundanese instances, and (2) manual writing by native speakers based on pre-defined topics, producing roughly 529 Javanese instances and 595 Sundanese instanecs. Each strategy undergoes rigorous quality control to ensure accuracy and reliability. In total, we create a high-quality test set of 2,124 instances across both languages.
To ensure the authenticity and quality of the dataset, we recruited 4 expert workers (2 per language) who are not only native Indonesian speakers but also fluent in Javanese and Sundanese. Each expert worker has a deep understanding of their respective language, culture, and customs. They have at least 10 years of experience speaking Javanese or Sundanese and possess strong linguistic and cultural expertise. The recruited workers, aged between 21 and 35 years, hold bachelor’s degrees and were carefully selected for their proficiency in both language and cultural knowledge.
Machine translation with human verification
As shown in Figure 1, we translate 500 randomly selected samples from the IndoCloze test set into Javanese and Sundanese using Google Translate. [^4] Google Translate was accessed in September 2024. Similar to the construction of training set, we choose Google Translate for its strong translation performance in both Javanese and Sundanese. To ensure quality, we employ two native speakers for each language and implement a two-stage quality control process. In Stage 1, the first worker manually corrects translation errors and localizes content by replacing entities (e.g., names, buildings, food) with culturally relevant alternatives. In Stage 2, a second worker validates the revised text and directly corrects any remaining errors from the first stage. This verification step ensures the accuracy and naturalness of the machine-translated test set. From this point forward, we refer to this data as MTtest.
Human-written Dataset
Each expert worker in Sundanese and Javanese is tasked with writing 600 short stories following the IndoCloze format: a four-sentence premise, a correct fifth sentence, and an incorrect fifth sentence. Stories are written based on 12 predefined topics, adhering to the same topic taxonomy used for training (see Figure 6 and Section 3.2). Each topic requires 50 stories, resulting in a total of 600 stories per language. [^5] See Appendix K for further details on the writing guidelines.
To ensure quality, each expert worker reviewed their peer’s written stories using the same pairing system as in the IndoCloze portion. The reviewing worker was presented with a premise and two randomized alternate endings from another worker’s story and was asked to identify the correct one. Instances incorrectly identified by the second worker were discarded, as they likely contained incorrect endings or exhibited ambiguity. After quality control, 529 Javanese and 595 Sundanese instances remained from the original 600 per language. From this point forward, we refer to this human-written dataset as HW.
4 Data Analysis
4.1 Overall Statistics
For LLM-assisted dataset creation, GPT-4o and Claude model demonstrated the highest efficiency, generating nearly 1,000 clean samples with minimal discarded output, while Mixtral was the least efficient, requiring significantly more samples to reach the same threshold (see Appendix F). The LLM-generated data does not tend to have a uniform topic distribution due to variations in broken samples.
In total, the LLM-generated datasets contain 72K sentences and approximately 557K words. The word distribution across sentence positions remains consistent across the six LLMs, with word counts per position relatively uniform and a median sentence length ranging from 5 to 10 words (see Appendix E).
The MT and Human-written datasets exhibit a similar word distribution pattern to the LLM-generated datasets. MT contains around 6K sentences with 44,5K words, while Human-written data has 6,7K sentences with 58,2K words. Despite the slight difference in word count, both datasets maintain a consistent distribution, with a median sentence length ranging from 4 to 11 words.
Model Sundanese Javanese
Coherence Fluency Correctness Cultural Rel. Coherence Fluency Correctness Cultural Rel.
GPT-4o 4.7 4.2 80 96 4.9 4.5 96 68
Claude 4.7 4.4 86 92 4.9 4.3 92 74
Cohere 3.4 3.0 28 46 4.5 4.1 70 52
Llama3.1 3.7 3.4 56 70 4.5 4.3 54 44
Gemma2 3.9 3.1 42 78 4.8 3.6 76 54
Mixtral 2.0 1.5 0 4 2.2 2.1 0 6
MTtest 5.0 5.0 100 14 5.0 5.0 100 18
HW 5.0 5.0 100 96 5.0 5.0 100 66
Table 2: Quality analysis of various models on Sundanese and Javanese. Higher scores indicate better performance in each category. “HW” denotes human-written dataset
Model Sundanese Javanese
#data #vocab LW (%) $\downarrow$ MATTR $\uparrow$ #data #vocab LW (%) $\downarrow$ MATTR $\uparrow$
GPT-4o 1000 3444 0 0.82 1000 3073 0 0.80
Claude 1000 3898 0 0.80 1000 3104 0 0.79
Cohere 1000 2654 3 0.65 1000 2254 3 0.68
Llama3.1 1000 3126 0 0.69 1000 2836 2 0.69
Gemma2 1000 3758 5 0.72 1000 3334 4 0.71
Mixtral 1000 4584 16 0.71 1000 4215 13 0.67
MTtest 500 3877 0 0.82 500 3620 3 0.81
HW 594 4693 0 0.84 529 3497 0 0.80
Table 3: Lexical diversity analysis of different models on Sundanese and Javanese.
4.2 Quality Analysis based on Human Evaluation
We evaluate the quality of LLM-generated, machine-translated, and human-written datasets through human assessment based on four key criteria: coherence, fluency, correctness, and cultural relevance. Specifically, we randomly select 50 samples (5–10% of the total dataset) for both Sundanese and Javanese and engage two native speakers for evaluation. Coherence and fluency are rated on a Likert scale from 0 to 5, while correctness and cultural relevance are assessed using binary annotation and reported as percentages. More details on the annotation guideline can be found in Appendix J. Table 2 summarizes the results of human evaluation, with scores for coherence and fluency being averaged between the annotators, while we count the percentage of data perceived as correct or culturally relevant by both annotators to aggregate scores on correctness and cultural relevance. The Pearson correlation between annotators for coherence and fluency ranges from 0.6 to 0.7, while Kappa scores for correctness and cultural relevance range from 0.4 to 0.6.
We observe that across LLMs, GPT-4o and Claude model consistently demonstated strong performance across all metrics, especially in correctness and cultural relevance. Notably, in terms of cultural relevance, both models achieved cultural relevance scores comparable to human written text, making them the best-performing LLMs in this aspect. This also indicates that GPT-4o and Claude model are capable of generating culturally appropriate content at a level similar to human writing. However the major issue is in correctness, where GPT-4o has 20% error in Sundanese. Machine translation with post-edit might produce a good coherence, fluency, and correctness. However, the data has lower cultural relevance score. This highlight that traditional approach with machine translation is not the best practice in low-resource
4.3 Lexical Diversity
We analyze the lexical diversity of LLM-generated data for Sundanese and Javanese using key metrics such as vocabulary size, moving-average type-token ratio (MATTR) Covington and McFall (2010), and the proportion of loanwords. To measure loanword presence, we manually review the top 100 most frequent words for each model.
As shown in Table 3, GPT-4o achieves a MATTR score of 0.82 for Sundanese and 0.80 for Javanese, making it highly comparable to human-written data. Claude also demonstrates strong lexical diversity, maintaining a rich vocabulary while avoiding borrowed words. Machine translation (MTtest) has a competitive MATTR score but contains 3% loanwords in Javanese.
Among the LLM-generated datasets, Mixtral has the largest vocabulary size but also introduced the highest proportion of loanwords, with 16% in Sundanese and 13% in Javanese, suggesting a significant reliance on non-native terms. In contrast, GPT-4o and Claude generate text entirely in Sundanese and Javanese without incorporating foreign words, highlighting their ability to produce better datasets.
In name generation, LLMs frequently produce common Javanese names as Ayu, Dwi, Bayu, Eko, and Sari, while Mixtral and Gemma2 showed greater diversity rather than other LLM. For Sundanese, LLMs primarily generate names like Lia, Budi, Dewi, Rina, whereas Llama3.1 and Gemma2 introduce more varied name choices. All these names are widely used in Indonesia. However, variations in honorific terms are minimal across all models, with only Pak and Bu (along with variants like Ibu and Bapak) appearing consistently.
5 Experiments and Analysis
5.1 Classification
Setup
We adopt the classification accuracy metric, as proposed in both Mostafazadeh et al. (2016) and Koto et al. (2022a), defined as the ratio of correctly classified cases to the total number of test cases, $\frac{correct}{testcases}$. We fine-tune all models using the training set in Javanese and Sundanese combined, and average the results over three runs to ensure robustness. In this study, we train several models and evaluate their adaptability in the two languages. We employ several models for our experiments: BiLSTM, and XLM-R (Conneau et al., 2019) (Details in Appendix A).
Training #Data BiLSTM XLM-R
ID 1K 47.7 ± 1.1 62.8 ± 0.4
MTTrain 2K 58.2 ± 0.6 69.4 ± 0.5
GPT-4o 2K 62.8 ± 0.6 72.7 ± 0.5
Claude 2K 64.9 ± 0.6 72.0 ± 0.8
Cohere 2K 57.1 ± 0.1 68.6 ± 0.8
Llama3.1 2K 60.1 ± 0.6 72.0 ± 0.4
Mixtral 2K 60.2 ± 0.5 54.9 ± 0.5
Gemma2 2K 55.3 ± 0.9 56.6 ± 0.7
All LLMs 12K 65.1 ± 0.1 72.5 ± 0.4
All LLMs + Filtering 1.2K 62.7 ± 0.3 70.4 ± 0.2
All LLMs + Filtering 0.6K 61.6 ± 0.7 69.4 ± 0.1
Table 4: Accuracy across different training sets. “All LLMs + Filtering” has two variants: (1) a curated set of 1,220 samples selected through the automatic filtering pipeline and (2) a subset of 674 samples containing only filtered data generated by GPT-4o and Llama3.1.
Overall Performance
Table 4 shows that training on LLM-generated datasets mostly improves performance over the baseline ID dataset, highlighting the decent quality of synthetic data for both languages. Among the LLM-generated training sets, Claude and GPT-4o perform particularly well, 3–4 points higher than machine translation (MTTrain). We also observe that data generated by Mixtral and Gemma2 yields worse accuracy than MTTrain. The full results of the classification experiments are shown in Appendix I
Train with Filtered Data
In As shown in Table 4, when using the full LLM-generated dataset (12K samples), XLM-R achieves an accuracy of 72.5, which is competitive with GPT-4o. After applying filtering, the training set is reduced to 1,220 and 674 samples (61% and 33.7% of the total training data generated by GPT-4o). Despite this substantial reduction in data size, accuracy drops by only 1–3%, highlighting the efficiency of training with a smaller yet more representative subset.
[…]
Accuracies on Machine-translated and Human-written Test Set
Figure 2 shows that when XLM-R is trained on MTTrain, it achieves similar accuracy (70%) on both MTTest and HW. Interestingly, compared to LLM-generated training data, MTTrain performs comparably or even better on MTTest but worse on HW. This is likely because MTTrain and MTTest originate from Koto et al. (2022b), resulting in a similar topical distribution. In contrast, the human-written test set and the LLM-generated training data were created using the same predefined topics.
Figure 3: Classification accuracy across topics.
Performance across Diferrent Topics
Figure 3 shows the performance of XLM-R in different topics. The chart shows that the data generated by GPT-4o and Claude consistently achieve high accuracy in most categories, but they also still underperform in some. The two models exhibit strong accuracy in categories such as Art, Socio-religious aspects of life, Agriculture and Wedding, and are generally weaker in categories like Fisheries and Trade, Death, Daily Activities and Pregnancy and Kids. It can also be seen that the two models give almost identical performance in Sundanese whereas GPT-4o seems slightly better in Javanese.
5.2 Generation
Setup
We use a four-sentence premise as input and train Indonesian LLM, Sahabat.AI (GoToCompany, 2024), to generate the fifth sentence in both Javanese and Sundanese. We use QLoRA (Dettmers et al., 2023) to perform supervised fine-tuning. The hyperparameters for training can be seen on Appendix A. Automatic evaluation methods are used to evaluate the model performance: ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2020).
Overall Performance
Table 5 presents the automated evaluation metrics for different training sets across Sundanese and Javanese. Training on LLM-generated datasets generally improves performance over the baseline ID dataset. However, MT achieves the highest scores in Sundanese for ROUGE-L and METEOR and outperforms all LLM-generated datasets across all metrics in Javanese. Among LLM-generated datasets, Llama3.1 performs best, particularly in Javanese, where it achieves the highest ROUGE-L and METEOR scores.
Train Set Sundanese Javanese
R-L M BS R-L M BS
ID 17.8 12.0 70.7 19.9 13.7 72.1
MT 19.1 13.4 70.4 21.8 16.7 72.5
Claude 18.1 12.7 71.6 20.4 15.5 72.2
Llama3.1 18.6 12.7 70.5 23.2 18.0 72.6
Table 5: Generation performance across different training sets for Sundanese and Javanese test sets. Metrics: ROUGE-L (R-L), METEOR (M), and BERTScore (BS).
Figure 4: Human evaluation on generation task. LLM means model trained with LLM-generated data
Human Evaluation
We conducted a human evaluation to compare models trained on machine-translated data with those trained on LLM-generated data, selecting Claude as the representative LLM model. Workers were presented with two randomized responses and asked to determine whether the first response was better, the second was better, or if both were equally good (draw). The results in Figure 4 indicate a slight preference for the model trained on LLM-generated data over MTTrain in both languages.
6 Conclusion
We explored the potential and limitations of LLM-generated data for commonsense reasoning and story generation in Javanese and Sundanese, introducing the first cloze dataset for these languages with high-quality test sets. Our preliminary analysis in classification and generation settings shows that GPT-4o and Claude-3 Opus demonstrate strong capabilities in generating plausible short stories but face challenges in fluency and cultural accuracy. Despite these limitations, our findings suggest that LLM-assisted data generation is a practical and effective approach for constructing datasets in low-resource languages.
7 Limitations
Cultural Nuance, Language, and Dialect Coverage This study acknowledges inherent limitations in the scope of its cultural nuance, language, and dialect coverage. While we carefully curated our data and covered 12 distinct cultural topics, our data sources and predefined topics may not fully capture the comprehensive range of cultural expressions within these multicultural languages, encompassing all dialects and regional variations. Furthermore, while we performed manual human evaluation to assess the quality of our generated data to some extent, the evaluation was limited in scale and depth. Future research should explore a broader array of cultural themes, incorporate diverse data sources, improve human evaluation methods, and deepen the dialectal representation to better reflect the diversity of the targeted languages. We also suggest future studies extends the research to other local languages in Indonesia.
Prompt and Hyperparameter Variation Due to resource constraints and the primary focus on data generation strategies, this research explored a limited range of LLM prompts and hyperparameter configurations. While prompt engineering and model training were performed with due diligence, it is important to acknowledge that alternative prompts or hyperparameter settings could potentially yield different results. Future research should investigate a wider spectrum of prompts and hyperparameters to identify configurations that maximize the performance of LLMs in generating culturally nuanced common sense reasoning in Javanese and Sundanese.
8 Ethical Considerations
All human-written datasets have been manually validated to ensure that harmful or offensive questions are not present in the dataset. We paid our expert workers fairly, based on the monthly minimum wage in Indonesia [^6] The average monthly minimum wage in Indonesia is approximately 3,000,000 IDR. The workload to complete all the tasks equates to roughly 8 days of full-time work. Each worker was paid 1,250,000 IDR accordingly. All workers were informed that their stories submitted would be used and distributed for research. Furthermore, no sensitive or personal information about the workers would be disclosed.
Appendix A Training Configurations
A.1 Classification
  • n-gram overlap: For this approach, we select the candidate with the highest overlap of n gram, measured by ROUGE-1 (Lin, 2004), calculated between the premise and the end. This metric ensures optimal alignment between the candidate and the reference text.
  • FastText-similarity: We used a similarity-based classification method, selecting the candidate with the highest cosine similarity between the premise and the ending, using the 300-dimensional FastText word embeddings trained in respective local languages (Mikolov et al., 2018). [^7] https://fasttext.cc/docs/en/crawl-vectors.html
  • Hierarchical BiLSTM: We used a two-level 200-dimensional BiLSTM architecture similar to Koto et al. (2022a), with 300-dimensional FastText representations of the respective language as input.
  • Pre-trained Language Models (PLM): We fine-tuned XLM-R (Conneau et al., 2019), a pre-trained multilingual language model.
For the LSTM model, we limited each sentence to a maximum of 30 tokens and trained it for up to 100 epochs, using early stopping with patience of 20 epochs. The model was trained with a batch size of 20, using the Adam optimizer and a learning rate of 0.1.
We set the maximum token length for the pre-trained language model to 450 for the premise and 50 for the ending sentence. The model was trained over 20 epochs with early stopping (patience set to 5), using a batch size of 40, Adam optimizer, an initial learning rate of 5e-6 for XLM-R, and a warm-up phase comprising 10% of the total training steps.
A.2 Generation
We do supervised fine-tuning on Gemma2-9B-CPT-Sahabat-AI-v1-base (GoToCompany, 2024) base model using QLoRA. The model was trained with 4 bit quantization, lora rank 64, lora alpha 128, using a batch size of 8, gradient accumulation of 8, learning rate 2e-4, with 1 epoch. We use Unsloth.ai to perform the training (Daniel Han and team, 2023).
A.3 Data Filtering Classifier
We train XLM-R model with the maximum token length of 1024 for the premise and 128 for the ending sentence. The model was trained over 26 epochs, using a batch size of 16, Adam optimizer, an initial learning rate of 1e-5 for XLM-R, and a warm-up phase comprising 5% of the total training steps.
Appendix B Prompt Template
Figure 10 shows the prompt template used for in-context learning, guiding the LLM to generate new Javanese and Sundanese.
Appendix C Filtered LLM-Generated Train Data Proportion
Initially, training data was generated using six different LLMs, with each model contributing approximately 16.67% of the total 12,000 samples (around 2,000 samples per model). However, after filtering the bad examples, the final dataset composition shifted. The data consist of 1,220 samples, the distribution is as follows: Claude (37.7%), GPT-4o (29.0%), LLama (8.6%), Cohere (4.3%), Mixtral (2.7%), and Gemma-2 (1.7%).
Appendix D Topic Distribution
Figure 5 illustrates the balanced distribution of topics in various datasets generated by different LLM (Claude, Cohere, GPT-4o, Mixtral, Llama3.1 and Gemma2). Each topic contains approximately 166 samples, with minor variations, ensuring an even distribution across all categories. Meanwhile, Figure 6 visualizes the human-written dataset, where topics are similarly well distributed across both local languages.
Figure 5: Topic Distribution Across Datasets
Figure 6: Topic Distribution in Human-Written Data
Appendix E Sentence Length Distribution
Figure 8 illustrates the distribution of sentence lengths across six datasets generated by various LLMs. Each dataset exhibits a similar distribution, with most sentence lengths ranging from 5 to 10 words. While Figure 7 shows the distribution of sentence lengths from MT and Human-written, and both dataset exhibit a similar pattern of distribution as LLM-generated data, with most sentence lengths are ranging from 4 to 11.
Figure 7: Sentence Length Distribution Across Datasets: The figure shows the distribution of sentence lengths across MT and Human-written. Each bar represents the frequency of sentence lengths for different sentence types in each dataset.
Figure 8: Sentence Length Distribution Across Datasets: The figure shows the distribution of sentence lengths across various datasets generated by different LLMs. Each bar represents the frequency of sentence lengths for different sentence types in each dataset.
Appendix F Overgeneration Distribution
Figure 9 shown, the number of generated samples needed to produce 1,000 clean training data for Javanese and Sundanese varies between models, where lower numbers indicate higher efficiency due to fewer unusable examples such as duplicates or repetitions of premises.
Figure 9: Samples required to generate 1,000 clean training data: lower values indicate higher efficiency.
Appendix G Top-5 Bigrams Trigrams
Table 6 presents the top-5 bigrams and trigrams found in the dataset, highlighting frequent patterns such as "Pak Budi" and "warga desa" (villagers), reflecting common themes in LLM-generated Javanese and Sundanese texts.
Bigram (unique: 103,022) Freq (%)
pak budi (Mr. Budi) 0.2488
seneng banget (really happy) 0.2136
warga desa (villagers) 0.1660
lan kanca-kancane (and their friends) 0.1406
dina acara (event day) 0.1077
Trigram (unique: 167,659) Freq (%)
seneng banget karo (very happy with) 0.670
warga desa padha (villagers will) 0.440
dina budaya sunda (in Sundanese culture) 0.417
nu aya di (the one in) 0.335
kabeh warga desa (every villager) 0.326
Table 6: Top bigram and trigram occurrences in the dataset.
Appendix H Agreement Score
We calculated the fluency and coherency agreement scores using Pearson’s correlation and computed the correctness and cultural relevance scores using Cohen’s kappa. To evaluate the quality of LLM-generated examples, the workers followed a predefined scoring guideline (Appendix J). The agreement scores for Javanese and Sundanese are summarized in Table 7.
Metric Javanese Sundanese
Fluency 0.74 0.76
Coherency 0.78 0.71
Correctness 0.73 0.50
Cultural Relevance 0.23 0.49
Table 7: Agreement Scores for Javanese and Sundanese
Appendix I Classification Test Performance on Javanese and Sundanese
Tables 8 and 9 present test performance on Sundanese and Javanese across different training datasets. XLM-R and BiLSTM outperform the n-gram and FastText baselines, with LLM-generated datasets achieving comparable or better results than machine translation. Notably, All $LLMs+Filtered$ training sets yield the best performance, especially on human-written test sets, highlighting the benefit of curated synthetic data.
Appendix J Workers Guidelines
Rubric for Story Evaluation
  • Fluency: Each sentence should be grammatically correct and fluent.
    • Score Range: 0 to 5
    • 5: All sentences are grammatically correct and very fluent.
    • 0: Sentences are grammatically incorrect and lack fluency.
  • Coherency: The story should be coherent, with all sentences logically connected.
    • Score Range: 0 to 5
    • 5: Story is highly coherent, with clear and logical flow between sentences.
    • 0: Sentences are disconnected and lack a logical sequence.
  • Correctness: The correct story closure should be valid, while the incorrect closure should clearly be wrong.
    • Binary Score:
    • 1: The correct ending is indeed correct, and the incorrect ending is clearly wrong.
    • 0: Either the correct ending is not valid, or the incorrect ending is not clearly wrong.
  • Cultural Relevance: The story should reflect appropriate cultural norms, values, or symbols relevant to the language.
    • Binary Score:
    • 1: The story contains relevant cultural norms/values/symbols for the corresponding language.
    • 0: The story lacks cultural relevance or includes irrelevant cultural aspects.
Scoring Guidelines:
  • Coherency: Ranges from 0 to 5, where 5 means each sentence is strongly connected and flows well with the previous and next sentence.
  • Fluency: Ranges from 0 to 5, where 5 indicates all sentences are grammatically sound and highly fluent.
  • Correctness: A binary score of 0 or 1 to ensure that the correct ending is truly valid and the incorrect ending is clearly wrong.
  • Cultural Relevance: A binary score of 0 or 1 to ensure the whole story contains appropriate and relevant cultural symbols and norms for the language being used.
Appendix K Topic Taxonomy and Story-Writing Guidelines
Topic Taxonomy
As part of the Javanese and Sundanese Cloze Project, a total of 300 stories must be created. These stories are evenly distributed across 12 predefined topic categories, with 25 stories per topic. Each story must reflect traditional Javanese and Sundanese values and customs, with attention to detail and coherence in the narrative. The topic categories and their subcategories are as follows:
  • Food: Breakfast, Lunch, Dinner, Snacks, Food souvenirs, Traditional foods and beverages, Eating habits, Cutlery, Cooking ware, Fruits.
  • Wedding: Traditions before marriage, Traditions during marriage, Traditions after marriage, Men’s wedding clothes, Women’s wedding clothes, Invited guests, Wedding location, Foods at a wedding, Gifts brought to a wedding.
  • Family Relationship: Relationships within the immediate family, Relationships in the extended family, Relations with society or neighbors, Clan or descendant system.
  • Pregnancy and Kids: Traditions during pregnancy, Traditions after birth, Caring for a newborn baby, Caring for toddlers, Caring for children, Caring for teenagers, Parent-child interactions as adults.
  • Death: When death occurs, The process of caring for a corpse, Traditions after burial, Mourners’ clothes, Inheritance matters.
  • Religious Holiday: Traditions before religious holidays, Traditions during religious holidays, Traditions after religious holidays.
  • Agriculture: What to plant, Traditions during planting, Harvest.
  • Fisheries and Trade: Traditions of taking care of livestock or fish, Buying and selling traditions.
  • Art: Musical instruments, Folk songs, Traditional dances, Use of art at specific events, Poetry or similar literature.
  • Traditional Games: Types of games, Locations where games are played.
  • Daily Activities: Morning activities, Afternoon activities, Evening activities, Leisure activities, Household tasks, Transportation.
  • Socio-religious Aspects of Life: Regular religious activities, Mystical elements, Traditional ceremonies, Lifestyle, Self-care, Traditional medicine, Traditional sayings.
Story-Writing Guidelines
Each story must consist of 4 sentences and two endings: one correct and one incorrect. The following example demonstrates how stories should be written (Table 10):
Ensure that all stories adhere to the topics and categories outlined in the taxonomy and reflect traditional values and cultural relevance.
Figure 10: Prompt template instructing LLM to generate a new example for Javanese and Sundanese, given a set of in-context examples.
Train Set Machine Translated - Test Human-written
n-gram FastText BiLSTM XLM-R n-gram FastText BiLSTM XLM-R
ID 54.6 63.0 51.4 ± 1.1 58.4 ± 0.6 55.7 62.5 49.3 ± 0.2 68.0 ± 0.6
MT 59.3 ± 0.2 67.2 ± 0.2 57.5 ± 0.1 74.3 ± 0.1
GPT-4o 63.5 ± 0.1 67.2 ± 0.1 62.8 ± 0.1 76.6 ± 0.1
Claude 63.5 ± 0.1 68.0 ± 0.1 63.7 ± 0.1 76.9 ± 0.1
Cohere 57.0 ± 0.1 62.6 ± 0.1 54.3 ± 0.1 70.5 ± 0.2
Llama3.1 60.4 ± 0.2 67.2 ± 0.1 58.1 ± 0.1 73.7 ± 0.1
Mixtral 57.4 ± 0.1 54.4 ± 0.1 59.2 ± 0.2 54.7 ± 0.1
Gemma2 59.4 ± 0.1 55.6 ± 0.1 61.2 ± 0.1 54.3 ± 0.1
All LLMs 63.4 ± 0.1 68.3 ± 0.1 63.1 ± 0.2 75.3 ± 0.1
All LLMS + Filtering 62.0 ± 0.1 67.2 ± 0.1 61.7 ± 0.1 71.5 ± 0.1
Table 8: Test Performance on Sundanese across Different Models and Training Data
Train Set Machine Translated - Test Human-written
n-gram FastText BiLSTM XLM-R n-gram FastText BiLSTM XLM-R
ID 63.5 61.0 49.7 ± 0.1 63.2 ± 0.7 72.4 62.9 40.6 ± 0.1 61.8 ± 0.1
MT 60.4 ± 0.1 70.5 ± 0.3 55.6 ± 0.2 65.7 ± 0.2
GPT-4o 56.7 ± 0.3 63.6 ± 0.1 71.0 ± 0.2 83.7 ± 0.1
Claude 58.9 ± 0.1 63.8 ± 0.1 73.4 ± 0.1 79.5 ± 0.1
Cohere 52.8 ± 0.2 64.2 ± 0.1 64.2 ± 0.1 77.2 ± 0.1
Llama3.1 58.5 ± 0.2 65.3 ± 0.1 63.5 ± 0.1 81.9 ± 0.1
Mixtral 58.1 ± 0.1 56.2 ± 0.1 66.2 ± 0.1 54.5 ± 0.1
Gemma2 52.4 ± 0.2 56.6 ± 0.1 48.3 ± 0.2 59.9 ± 0.1
All LLMs 60.4 ± 0.3 66.0 ± 0.1 73.4 ± 0.1 80.6 ± 0.1
All LLMs + Filtering 53.9 ± 0.1 63.6 ± 0.1 73.0 ± 0.1 79.5 ± 0.1
Table 9: Test Performance on Javanese across Different Models and Training Data
Topic Category Javanese Translated
Javanese
Food Traditional foods and beverages Premise: Laras seneng banget karo wedang uwuh, minuman tradisional saka Yogyakarta. Saben sore, dheweke mesthi ngombe wedang uwuh kanggo njaga kesehatan lan semangat. Wedang uwuh digawe saka campuran rempah-rempah alami sing wangi. Laras percaya yen wedang uwuh bisa ngusir rasa kesel lan lemes. Correct: Saben sore, dheweke seneng bareng-bareng karo kanca-kanca ngombe wedang uwuh. Incorrect: Dheweke ngindhari ngombe wedang uwuh saben sore. Premise: Laras is really happy with wedang uwuh, a traditional drink from Yogyakarta. Every evening, she must always drinks wedang uwuh to maintain her health and vitality. Wedang uwuh is made from a mixture of fragrant natural spices. Laras believes that wedang uwuh can drive away fatigue and weakness. Correct: Every evening, she enjoys drinking wedang uwuh with her friends. Incorrect: She avoids drinking wedang uwuh every evening.
Sundanese
Wedding Traditions when getting married Premise: Kulawargi bapak Syakur nembe pisan hajat nikahan. Anjeunna parantos nikahkeun putrana anu bungsu, Anton. Acara sawéranna heboh pisan, aya rupi-rupi doorprize. Abi emut aya hiji ibu-ibu nu kéngéng rice cooker. Correct: Aya ibu-ibu nu kéngéng doorprize tina sawéran hajatan pak Syakur. Incorrect: Pak Syakur masihan rice cooker kumargi ibu-ibu eta orang Premise: My family recently had a wedding. He has married his youngest son, Anton. The event was very exciting, there were various door prizes. I remember there was a mother who received a rice cooker. Correct: There were mothers who received door prizes from Mr. Syakur’s wedding. Incorrect: Mr. Syakur gave the rice cooker because the mothers were improtant people.
Table 10: Example of a story with correct and incorrect endings.